OpenStudy (anonymous):

Integration: \[\int {\frac{{dx}} {{{x^4} + 1}}} \]

OpenStudy (anonymous):

I have done to this \[\frac{1} {{{x^4} + 1}} = \frac{{\frac{x} {{2\sqrt 2 }} + \frac{1} {2}}} {{{x^2} + \sqrt 2 x + 1}} + \frac{{\frac{{ - x}} {{2\sqrt 2 }} + \frac{1} {2}}} {{{x^2} - \sqrt 2 x + 1}}\]

OpenStudy (anonymous):

And I don't know how to do next. ....

OpenStudy (anonymous):

what about this pls help t=3n+6n

OpenStudy (anonymous):

t=9n lolz

OpenStudy (anonymous):

hmm Okay I have this divide numerator and denominator by \(x^2\)

OpenStudy (anonymous):

\[\int \frac{\frac{1}{x^2}}{x^2+ \frac{1}{x^2} +2 -2}\]

OpenStudy (anonymous):

pls help need answer t=6-3n .....11 years old don't understand homework lol

OpenStudy (lalaly):

Note that\[ x^4 + 1 = (x^2 + √2 x + 1)(x^2 -√2x + 1).\] Now break \[\frac{1}{[ (x^2 + √2 x + 1)(x^2 -√2x + 1)]}\] into partial fractions and integrate.